Use elimination to solve the system of equations.
-2x+y+3z=20
-3x+2y+z=21
3x-2y+3z=-9

Answer by checkley77(12844)

-2x+y+3z=20
-3x+2y+z=21
3x-2y+3z=-9 ADD THESE 2 EQUATIONS.
----------------
4Z=12
Z=12/4
Z=3 ANS.
-2X+Y+3*3=20
-2X+Y+9=20 NOW MULTIPLY THIS EQUATION BY -2 & ADD THEM.
-3X+2Y+3=21
4X-2Y-18=-40
--------------
X-15=-19
X=-19+15
X=-4 ANS.
3*-4-2Y+3*3=-9
-12-2Y+9=-9
-2Y=-9+12-9
-2Y=-6
Y=-6/-2
Y=3 ANS.